======20240205========== =====Time===== =====Attendance===== PICCSAR: Ian, Katie Obs staff: Onkar, Tara, Roy =====Spaces Occupied===== 401,433 =====Goals===== take data for PICCSAR =====Frame Table===== ^ Time Observed ^ Source Name ^ Filter ^ Int Time [s] ^ Mag (filter) ^ RA [h] ^ DEC [deg] ^ Source Type ^Scope ^ Purpose ^ Notes^ | 202402052100 | HD 432 | V |12ms | 1.25(V) | 00.09 | 58.04 | Star | Main | piccsarr |15.042 w/ ag| | 202402052105 | HD 7927 | V |12ms | 1.25(V) | 20 | 45 | Star | Main | piccsarr |15.042 w/ ag| | 202402052110 | HD 61421(procyon)| V |12ms | 1.25(V) | 20 | 45 | Star | Main | piccsarr |15.042 w/ ag| | 202402052110 | HD 21291| V |12ms | 1.25(V) | 20 | 45 | Star | Main | piccsarr |15.042 w/ ag| | 202402052110 | GB6 J0601+5315| V |12ms | 1.25(V) | 20 | 45 | blazar | Main | piccsarr |15.042 w/ ag| | 202402052220 | G99-37| V |30s | 1.25(V) | 20 | 45 | white dwarf | Main | piccsarr |15.042 w/ ag| | 202402052320 | MKN 421| V |30s | 1.25(V) | 20 | 45 | white dwarf | Main | piccsarr |15.042 w/ ag| | 202402052340 | PG 0934+554| V |30s | 1.25(V) | 20 | 45 | white dwarf | Main | piccsarr |15.042 w/ ag| | 202402052350 | MKN 421| V |30s | 1.25(V) | 20 | 45 | white dwarf | Main | piccsarr |15.042 w/ ag| | 202402050030 | G99-37| V |20s | 1.25(V) | 20 | 45 | white dwarf | Main | piccsarr |15.042 w/ ag| | 202402050030 | OJ 287| V |30s | 1.25(V) | 20 | 45 | white dwarf | Main | piccsarr |15.042 w/ ag| =====Log===== slew to HD 432:can't see anything do velcro maneuver still can't see anything \\ slew to Jupiter so so very out of focus slew to HD 432 focus currently: 2498.1 slew to HD 7927 slew to HD 61421 G99-37 also known as WD 0548−001 not in database so could not slew slew to HD 21291 \\ slew to GB6 J0601+5315 couldn't spot so temporarily changed int time to 1s still very dim \\ slew to M42 for fun \\ trying G99-37 again but 14.4 is beyond our ability \\ try and fail at Mkn 421 \\ switch to PICSI but GUI not being able to open for some reason, instead using lab laptop\\ broad focused by one too many significant figures seemed to stop DFM from responding\\ dome went a little wild, Onkar went up to realign the dome controls and computer\\ still could not see anything after the great unfocusing and switching the focus on the DFM was still not changing it\\ I went to the bathroom and it was fixed when I got back \\ slew to PG 0934+554\\ slew to MKN 421\\ slew to G99-37 again\\ last but not least slew to OJ 287\\ =====ToDo===== =====Notes===== =====Weather===== =====Data===== {{tag>[picsarr]}}